OpenStudy (anonymous):

PLEASE SOMEONE HELP!! WILL GIVE MEDAL AND FAN! Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

OpenStudy (anonymous):

@kyusakazaki

OpenStudy (anonymous):

@shkrina

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

write the question clearly

OpenStudy (anonymous):

it is clearly...... thats exactly how it is written @shkrina

OpenStudy (anonymous):

6x^5y^10

OpenStudy (anonymous):

part1 the production is commutative so we should regroupe numbers and x and y , then we have to calculate the sum of the exponential the calculations ok ,i can't do part one i'll do just the calculations ok ,(3x2x)*(3*2)(y^(6+4))=(6x^2)(6y^10)=36(x^2)(y^10)

OpenStudy (uri):

(3x3y4)(2x2y6) Mutiply the first bracket with the whole second one: 3x^3y^4*2x^2y^6 =6x^5y^10